OpenStudy (anonymous):

find the x-coordinate of all points on the curve y=36xsin(17x) + 306sqrt3(x^2) -33 domain pi/34 < x < 3pi/34 where the tangent line passes through the point (0, -33)(not on the curve)

OpenStudy (anonymous):

\[y = 36xsin(17x) + 306\sqrt{3}x^2 -33 \]

OpenStudy (anonymous):

\[\frac{ \Pi }{ 34 } < x < \frac{ 3\Pi }{ 34 }\] is the domain

OpenStudy (anonymous):

answer is x1 = 5Pi/102 and x2 = 7Pi/102 i need to know how to get it though